Alex's age is 3 year more than twice Vincent's age. in 4 years, the sum of their age is 58.

Answer by Maths68(1474)

Let
Present age of Alex = a
Present age of Vincent = v

Given
=======
Alex's age is 3 year more than twice Vincent's age
a=2v+3...........(1)
In 4 years
Alex will be = a+4
Vincent will be = v+4
Then
the sum of their age is 58
a+4+v+4=58
a+v+8=58
a+v=58-8
a+v=50.............(2)
Put the value of a from (1) to (2)
2v+3+v=50
3v=50-3
3v=47
3v/3=47/3
v=15.66
Put the value of v in (1)
a=2v+3...........(1)
a=2(15.66)+3
a=31.33+3
a=34.33



Present age of Alex = a = 34.33 years old (approximately 34 years 3 months 28 days)
Present age of Vincent = v = 15.66 years old (approximately 15 years 7 months 27 days)